OpenStudy (anonymous):

solve the following equation -4(2x-6)=2(15-5x)

OpenStudy (anonymous):

just upfold... -4*2x -4*(-6)-2*15-2*15x simplify and solve it for x

OpenStudy (anonymous):

-8x+24 = 30-10x. simplify to get 2x = 6. x= 3 hope this helps!

OpenStudy (anonymous):

correction...: -4*2x -4*(-6)=2*15-2*15x

OpenStudy (anonymous):

still lost :( inik

OpenStudy (anonymous):

do you understand what I did up there...?

OpenStudy (anonymous):

If you go back to the original post... It is totally worked out... I am telling you your multiple choice answers are all incorrect. the answer is x = 3

OpenStudy (anonymous):

lets try: -8x+24=30-10x add 10x to both sides: 2x+24=30 -24 both sides: 2x=6 x=6/2 x=3 - answer